### v2

# Bayesian BiLO: Bilevel Local Operator Learning for Efficient Uncertainty Quantification of Bayesian PDE Inverse Problems with Low-Rank Adaptation

Abstract
Uncertainty quantification in PDE inverse problems is essential in many applications. Scientific machine learning and AI enable data-driven learning of model components while preserving physical structure, and provide the scalability and adaptability needed for emerging imaging technologies and clinical insights. We develop a Bilevel Local Operator Learning framework for Bayesian inference in PDEs (B-BiLO). At the upper level, we sample parameters from the posterior via Hamiltonian Monte Carlo, while at the lower level we fine-tune a neural network via low-rank adaptation (LoRA) to approximate the solution operator locally. B-BiLO enables efficient gradient-based sampling without synthetic data or adjoint equations and avoids sampling in high-dimensional weight space, as in Bayesian neural networks, by optimizing weights deterministically. We analyze errors from approximate lower-level optimization and establish their impact on posterior accuracy. Numerical experiments across PDE models, including tumor growth, demonstrate that B-BiLO achieves accurate and efficient uncertainty quantification.
keywords:
Bayesian inference, PDE inverse problems, Bilevel Local Operator Learning, Low-Rank Adaptation, Hamiltonian Monte Carlo
pacs:
MSC Classification]65M32, 62F15, 68T07
1 Introduction
Uncertainty quantification in computational models governed by partial differential equations (PDEs) is essential across scientific and engineering disciplines, including seismic imaging [dengOpenFWILargeScaleMultiStructural2023, martinStochasticNewtonMCMC2012, yangSeismicWavePropagation2021, bui-thanhExtremescaleUQBayesian2012], personalized medicine [lipkovaPersonalizedRadiotherapyDesign2019, chaudhuriPredictiveDigitalTwin2023], climate modeling [senGlobalOptimizationMethods2013], and engineering design [cookOptimizationTurbulenceModel2019]. These fields frequently require solving PDE inverse problems within a Bayesian framework [stuartInverseProblemsBayesian2010], where the objective is to estimate the posterior distribution of model parameters given observed data. Typically, this process involves sampling PDE parameters and repeatedly solving the forward PDE. A related approach, PDE-constrained optimization, aims to identify a single optimal parameter estimate, thus offering computational efficiency but lacking the ability to quantify uncertainty. Beyond the cost and challenges of solving PDEs, mathematical models may be misspecified due to limited scientific understanding and therefore require continual refinement as new data become available. Scientific machine learning offers a framework for learning components of PDE models directly from data while retaining physical structure. Moreover, the rapid pace of new imaging technologies and clinical insights demands computational tools that are fast, scalable, and adaptable—capabilities well aligned with modern AI/ML methods. Such needs have driven growing interest in scientific machine learning methods that incorporate data-driven learning of model components.
The Bilevel Local Operator learning (BiLO) method has recently shown promise for solving PDE inverse problems in the constrained optimization framework [zhangBiLOBilevelLocal2025a]. A brief review of BiLO is provided in the supplementary material (SM). In this work, we introduce a novel framework, Bayesian BiLO (B-BiLO), which redesigns this bilevel structure for the more challenging setting of Bayesian inference. B-BiLO enables robust uncertainty quantification by integrating local operator learning with gradient-based MCMC sampling and transfer learning techniques for efficiency.
Addressing Bayesian PDE inverse problems involves two design choices: the PDE solver and the sampling method. PDE solutions can be computed using numerical methods (such as finite difference or finite element methods), physics-informed neural networks (PINNs)[raissiPhysicsinformedNeuralNetworks2019], or neural operators such as Fourier Neural Operator [liFourierNeuralOperator2021] and DeepONet [luLearningNonlinearOperators2021]. Each approach has its advantages and limitations. Numerical methods provide fast, accurate, and convergent solutions but can struggle with high-dimensional PDEs [grossmannCanPhysicsinformedNeural2024]. PINNs offer a mesh-free approach, effective in some higher-dimensional PDE problems, and a unified framework for inverse problems. Neural operators, while requiring many synthetic numerical solutions to train, enable rapid evaluations and serve as PDE surrogates; however, their accuracy typically degrades when evaluated outside the distribution of the training dataset [nelsenOperatorLearningMeets2025]. For sampling, Markov Chain Monte Carlo (MCMC) [hastingsMonteCarloSampling1970] methods can be either gradient-based (e.g., Hamiltonian Monte Carlo [nealMCMCUsingHamiltonian2011, nealBayesianLearningNeural1996, girolamiRiemannManifoldLangevin2011]) or derivative-free (e.g., Ensemble Kalman Sampling [garbuno-inigoInteractingLangevinDiffusions2020]). Gradient-based methods efficiently explore posterior distributions, but computing the gradient can be expensive (compared with gradient-free approaches) or impractical (e.g., when the PDE solver is a black box).
Numerical PDE solvers can be paired with gradient-based sampling methods. The Adjoint Method computes gradients of the likelihood function with respect to PDE parameters [bui-thanhSolvingLargescalePDEconstrained2014]. While the adjoint method ensures accurate solutions and gradients, deriving and regularizing the adjoint equations for complex PDEs can be challenging and solving both the forward and adjoint equations numerically can be computationally intensive. When only the forward solver is available, it is more practical to use derivative-free sampling methods such as Ensemble Kalman Sampling (EnKS), which employ ensembles of interacting particles to approximate posterior distributions [garbuno-inigoInteractingLangevinDiffusions2020]. Additionally, variants of Ensemble Kalman Inversion (EnKI) can efficiently estimate the first two moments in Bayesian settings [huangEfficientDerivativefreeBayesian2022, huangIteratedKalmanMethodology2022, pensoneaultEfficientBayesianPhysics2024].
Among neural PDE solvers, Physics-Informed Neural Networks (PINNs) [raissiPhysicsinformedNeuralNetworks2019] have become widely adopted. In PINNs, the PDE residual is enforced as a soft constraint. By minimizing a weighted sum of the data loss and the PDE residual loss, PINNs can fit the data, solve the PDE, and infer the parameters at the same time. For uncertainty quantification, Bayesian PINNs (BPINNs) [yangBPINNsBayesianPhysicsinformed2021] represent PDE solutions using Bayesian neural networks, and use Hamiltonian Monte Carlo (HMC) [duaneHybridMonteCarlo1987, nealBayesianLearningNeural1996] to sample the joint posterior distributions of PDE parameters and neural network weights. Derivative-free methods, such as EnKI, can also be effectively combined with neural PDE solvers, as demonstrated in PDE-constrained optimization problems [guthEnsembleKalmanFilter2020] and Bayesian PINNs [pensoneaultEfficientBayesianPhysics2024].
Neural Operators (NOs) approximate the PDE solution operator (parameter-to-solution map), providing efficient surrogate models for forward PDE solvers [kovachkiNeuralOperatorLearning2022]. Once trained, these surrogates integrate seamlessly with Bayesian inference frameworks, leveraging rapid neural network evaluations for posterior distribution sampling using both gradient-based and derivative-free methods [pathakFourCastNetGlobalDatadriven2022, luMultifidelityDeepNeural2022, maoPPDONetDeepOperator2023]. Prominent examples include the Fourier Neural Operator (FNO) [liFourierNeuralOperator2021, liPhysicsInformedNeuralOperator2024, whitePhysicsInformedNeuralOperators2023, akyildizEfficientPriorCalibration2025], Deep Operator Network (DeepONet) [luLearningNonlinearOperators2021, wangLearningSolutionOperator2021], and In-Context Operator (ICON) [yangIncontextOperatorLearning2023], among others [oleary-roseberryDerivativeInformedNeuralOperator2024, molinaroNeuralInverseOperators2023]. Nevertheless, neural operator performance heavily depends on the training dataset, potentially limiting accuracy in PDE solutions and consequently degrading posterior distribution fidelity [caoResidualbasedErrorCorrection2023].
Our framework integrates key practical advantages from existing sampling and PDE-solving methods. We employ gradient-based MCMC, which generally explores the posterior distribution more efficiently than derivative-free approaches. The PDE solution and the corresponding likelihood gradients with respect to parameters are computed using a neural solver within a novel bilevel optimization framework. This design is flexible across PDE models, and does not require either deriving the adjoint equations or curating pretraining datasets. Our methodology improves accuracy through strong enforcement of PDE constraints without balancing data fitting and PDE residuals as in PINNs, and enhances efficiency by optimizing network weights deterministically rather than sampling in a high-dimensional weight space as in Bayesian neural networks. The key contributions of this work are as follows:
  • We introduce a Bayesian Bilevel local operator (B-BiLO) learning approach that enforces strong PDE constraints, leading to improved accuracy in uncertainty quantification and parameter inference.
  • We avoid direct sampling in the high-dimensional space of Bayesian neural networks, leading to more efficient sampling of unknown PDE parameters or functions.
  • We apply Low-Rank Adaptation (LoRA)[huLoRALowRankAdaptation2021] to enhance both efficiency and speed in fine-tuning and sampling, significantly reducing computational cost while maintaining accuracy.
  • We prove there is a direct link between the tolerance for solving the lower level problem to approximate the local solution operator and the accuracy of the resulting uncertainty quantification. In particular, we estimate the error in the posterior distribution due to inexact minimization of the lower level problem.
2 Results
2.1 Method Overview
This section provides an overview of B-BILO. Full methodological details and theoretical analysis are provided in Section 4 and the Supplementary Material (SM).
PDE Bayesian Inverse Problem
We consider problems governed by partial differential equations (PDEs). Let $u:\Omega\to\mathbb{R}$ be a function (the state variable) defined on a domain $\Omega\subset\mathbb{R}^{d}$. The governing PDE, together with boundary and initial conditions, can be written compactly as
$\mathcal{F}[u,\theta](\mathbf{x})=\mathbf{0},\qquad\mathbf{x}\in\Omega,$
where $\theta\in\mathbb{R}^{m}$ are the PDE parameters and $\mathcal{F}$ is the associated differential operator, defined by
$\mathcal{F}[u,\theta](\mathbf{x}):=F(\mathcal{D}^{k}u(\mathbf{x}),\dots,\mathcal{D}u(\mathbf{x}),u(\mathbf{x}),\theta),\qquad\mathbf{x}\in\Omega,$ (1)
with $\mathcal{D}^{k}$ denoting the $k$-th order derivative operator. The solution to the PDE depends implicitly on $\theta$; we denote this solution by $u(\cdot,\theta)$.
Let $P(\theta)$ be the prior distribution of the PDE parameters. Let $D_{u}$ be the observed data, which is typically a noisy observation of the state $u$ at a set of measurement points. The likelihood $P(D_{u}|\theta)$ measures the probability of observing the data $D_{u}$ given the PDE parameters $\theta$. Since it is usually a functional of the state $u$, thus we also denote it as $\ell[u]$. For example, if we assume $D_{u}=\{\hat{u}_{i}\}_{i=1}^{N}$ with $\hat{u}_{i}=u(\mathbf{x}_{i},\theta)+\eta_{i}$, where $\eta_{i}$ is Gaussian noise of mean 0 and variance $\sigma_{d}^{2}$, then the likelihood of the observed data $D_{u}$ given the PDE parameters $\theta$ is
$\ell[u]=\prod_{i=1}^{N}\frac{1}{\sqrt{2\pi\sigma_{d}^{2}}}\exp\left(-\frac{1}{2\sigma_{d}^{2}}\left|u(\mathbf{x}_{i})-\hat{u}_{i}\right|^{2}\right)$ (2)
which is $P(D_{u}|\theta)$ if $u$ is the solution at $\theta$. Let $P(\theta|D_{u})$ be the posterior distribution of the PDE parameters. By Bayes’ theorem, $P(\theta|D_{u})\propto P(D_{u}|\theta)P(\theta)$. The potential energy, $\mathcal{U}[u,\theta]$, which depends on both the state $u$ and the PDE parameters $\theta$, is defined as the negative of the log-posterior distribution
$\mathcal{U}[u,\theta]=-\log P(\theta|D_{u})=-\log\ell[u]-\log P(\theta)$ (3)
We also denote $U(\theta)=\mathcal{U}[u(\cdot,\theta),\theta]$, as the potential energy ultimately depends on $\theta$ if the state $u$ is the solution at $\theta$. Sampling the posterior distribution with gradient-based MCMCs, such as Langevin Dynamics (LD) or Hamiltonian Monte Carlo (HMC), requires computing the gradient of the potential energy $\nabla_{\theta}U(\theta)$ to effectively explore the posterior distribution.
Bayesian Bilevel Local Operator Learning (B-BiLO).
We consider functions of the form $u(\mathbf{x},\theta)$ that depend on both the spatial-temporal variable $\mathbf{x}$ and the PDE parameters $\theta$. We call $u(\cdot,\theta)$ a local operator at $\theta$ if
$\mathcal{F}[u(\cdot,\theta),\theta](\mathbf{x})=\mathbf{0}\quad(\text{Condition 1}),$
$d_{\theta}\mathcal{F}[u(\cdot,\theta),\theta](\mathbf{x})=\mathbf{0}\quad(\text{Condition 2}),$
for all $x\in\Omega$, where $d_{\theta}$ is the total derivative with respect to $\theta$. Condition 1 requires the residual of the PDE to be zero, i.e., that $u(\cdot,\theta)$ is a solution of the PDE at $\theta$. Condition 2 requires the total derivative of the PDE residual with respect to $\theta$ to be zero, ensuring that the derivative of $u(\cdot,\theta)$ with respect to $\theta$ remains consistent with the PDE constraints. This ensures that small perturbations in $\theta$ lead to small perturbations in $u(\cdot,\theta)$ that still approximately satisfy the PDE constraints. In particular, if we consider a local perturbation $\theta\to\theta+\delta\theta$, a formal Taylor expansion of $\mathcal{F}$ about $\theta$ is zero to second order in $\delta\theta$ (e.g., see [zhangBiLOBilevelLocal2025a]).
In this work, we represent the local operator by a neural network $u(\mathbf{x},\theta;W)$, where $W$ are the weights of the neural network. The two conditions for $u$ to be a local operator at $\theta$ lead to the definition of the local operator loss $\mathcal{L}_{\text{LO}}$:
$\mathcal{L}_{\text{LO}}(\theta,W)=\norm{\mathcal{F}[u(\cdot,\theta;W),\theta]}^{2}_{2}+w_{\rm rgrad}\norm{d_{\theta}\mathcal{F}[u(\cdot,\theta;W),\theta]}^{2}_{2},$ (4)
We call the first term the residual loss and the second term the residual-gradient loss. Both terms can be approximated using the Mean Squared Error (MSE) at a set of collocation points. The weight $w_{\rm rgrad}>0$ is a scalar hyperparameter that controls the relative importance of the residual-gradient loss. In principle, both terms can be minimized to nearly zero. For fixed $\theta$, by minimizing the local operator loss $\mathcal{L}_{\text{LO}}(\theta,W)$ with respect to the weights $W$, we approximate the local operator $u(\mathbf{x},\theta)$ around $\theta$.
We consider the following Bayesian inference problem with a bilevel structure:
[…]
At the upper level, we sample the PDE parameters $\theta$ from the posterior distribution using Hamiltonian Monte Carlo (HMC) [nealMCMCUsingHamiltonian2011]. At the lower level, we train a network to approximate the local operator by minimizing the local operator loss with respect to the weights of the neural network.
Algorithm and Theoretical Analysis
We use HMC with the leapfrog integrator to sample the posterior distribution at the upper level. HMC requires computing the gradient of the potential energy $\nabla_{\theta}U(\theta)$ at various values of $\theta$, which require solving the lower-level problem to some tolerance $\epsilon$. Detailed algorithms are provided in Sec. 4.2. Under mild stability/smoothness assumptions, we show that the error in the gradient $\nabla_{\theta}U(\theta)$ and the error in the posterior distribution, both due to inexact minimization of the lower level problem, are $O(\epsilon)$. The proofs are provided in Sec. 4.3 and SM.
Advantages over existing methods
Many existing methods, such as Bayesian PINNs (BPINNs) [yangBPINNsBayesianPhysicsinformed2021, linkaBayesianPhysicsInformed2022], represent the PDE solution using a Bayesian neural network $v(\mathbf{x};W)$, (note that $\theta$ is not an input to the network), where $W$ follows a prior distribution $P(W)$. Uncertainty in $W$ induces uncertainty in the PDE residual $D_{f}=\{(\mathbf{x}_{f}^{(i)},\eta_{f})\}_{i=1}^{N_{f}}$, $\eta_{f}\sim\mathcal{N}(0,\sigma_{f}^{2})$, and give rise to a likelihood function for the residual term:
$P(D_{f}|W,\theta)=\prod_{i=1}^{N_{f}}\frac{1}{\sqrt{2\pi\sigma_{f}^{2}}}\exp\left(-\frac{1}{2\sigma_{f}^{2}}\left|\mathcal{F}[v(\mathbf{x}_{f}^{(i)};W),\theta]\right|^{2}\right)$ (6)
which serves as a soft constraint to enforce the PDE, and depends on $\sigma_{f}$. Assuming $\theta$ and $W$ are independent, the following joint posterior is sampled using HMC.
$P(W,\theta|D_{u},D_{f})\propto P(D_{u}|W)P(D_{f}|W,\theta)P(W)P(\theta).$ (7)
These would require sampling the challenging high-dimensional joint posterior of the PDE parameters $\theta$ and network weights $W$. Making $\sigma_{f}$ small enforces the PDE strongly, but also makes the loss ill-conditioned and leads to inefficient sampling. In contrast, B-BiLO avoids these issues by optimizing $W$ deterministically for each $\theta$, and thus only the low-dimensional parameter space $\theta$ is sampled. This approach removes the need for priors on weights, avoids the difficulties setting an artificial residual noise parameter, and enables more stable and efficient exploration of the posterior. A detailed review of BPINNs and it’s challenges are provided in SM.
Efficient Fine-Tuning via LoRA
To increase efficiency, we apply Low-Rank Adaptation (LoRA) [huLoRALowRankAdaptation2021] to reduce the number of trainable parameters at the lower level. Let a hidden-layer weight be $W\in\mathbb{R}^{p\times p}$. Across parameters $\theta$ and an initial guess $\theta_{0}$, BiLO updates only a low-rank increment
$W^{*}(\theta)\approx W^{*}(\theta_{0})+AB,\qquad A\in\mathbb{R}^{p\times r},\;B\in\mathbb{R}^{r\times p},\;r\ll p,$
while keeping the $W^{*}(\theta_{0})$ fixed. We refer to this procedure as “LoRA(rr)”, denoting LoRA fine-tuning with rank $r$ (Algorithm 3), whereas “Full FT” denotes full fine-tuning of all weights (Algorithm 2). Details of the neural network architecture and implementation of LoRA are provided in Methods. 4.1 and 4.4.
Schematic of B-BiLO with LoRA
A schematic of the B-BiLO framework is shown in Figure 1. The figure is based on a model boundary value problem $\theta u_{xx}=\sin(\pi x)$ with the boundary condition $u(0)=u(1)=0$. In this case, the full operator is given by $u(x,\theta)=\sin(\pi x)/(\pi^{2}\theta)$, which solves the PDE for all $\theta>0$. Figure 1 panel (A) illustrates the upper-level problem: the goal is to sample the PDE parameters $\theta$ from the posterior distribution. For gradient-based MCMC methods, we need to compute the gradient of the potential energy $\nabla_{\theta}U(\theta)$ at various values of $\theta$ (orange arrows). Therefore, at any fixed $\theta$, we need access to the local behavior of the potential energy, shown as the blue strip near $\theta_{1}$ and $\theta_{2}$.
In Figure 1 panel (B), the gray surface is the full PDE operator $u(x,\theta)$. At some particular $\theta_{1}$, the local operator $u(x,\theta_{1})$ approximates a small neighborhood of the full operator near $\theta_{1}$ (blue surface). The condition of the local operator is that the residual and the gradient of the residual with respect to $\theta$ are both zero at any fixed $\theta$. Knowing the PDE solution in the neighborhood of $\theta_{1}$ allows us to compute the gradient of the potential energy at $\theta_{1}$.
Figure 1 panel (C) illustrates the lower-level problem and LoRA: we train a neural network to approximate the local operator at different $\theta$, e.g. $\theta_{1}$ and $\theta_{2}$. The local operator at each $\theta_{i}$ is approximated by a neural network $u(x,\theta_{i};W)$ and is obtained by minimizing the local operator loss $\mathcal{L}_{\text{LO}}(\theta_{i},W)$. Instead of updating all the weights $W$ at each $\theta_{i}$, we use LoRA to update the weights for different $\theta_{i}$, which significantly reduces the number of trainable parameters and thus the computational cost.
Figure 1: A schematic for solving the Bayesian PDE inverse problem using BiLO and LoRA. (A) At the upper level, we sample the PDE parameters $\theta$ from the posterior distribution using gradient-based MCMC. (B) The full operator $u(x,\theta)$ (gray surface) and its local approximation $u(x,\theta_{1})$ and $u(x,\theta_{2})$ (blue surface), which satisfies vanishing residual and residual-gradient conditions. The local operator suffices to compute the gradient of the potential energy $U(\theta)$ at $\theta_{1}$ and $\theta_{2}$. (C) At the lower level, we train the neural network to approximate the local operator (blue surface) at different $\theta$. As $\theta$ changes, only low rank adaptation (LoRA) is used to update the neural network weights $W$.
2.2 Example 1: Nonlinear Poisson Equation
In this section, we compare B-BiLO with BPINN [yangBPINNsBayesianPhysicsinformed2021] on a nonlinear Poisson problem to demonstrate the advantages of B-BiLO: physical uncertainty quantification, stable sampling with larger step sizes, and reduced computational cost. We consider the 1D nonlinear Poisson equation [yangBPINNsBayesianPhysicsinformed2021]:
$\lambda u_{xx}(x)+k\tanh(u(x))=f(x),\quad x\in(-0.7,0.7)$ (8)
where $\lambda=0.01$, and both $f$ and the boundary conditions are derived from the ground-truth solution $u^{\rm GT}(x)=\sin^{3}(6x)$ with $k^{\rm GT}=0.7$. The goal is to infer $k$ from observations of $u$, with the prior that $k$ is uniformly distributed on $[0.4,1]$. A key feature of this problem is that the solutions are odd: both $u$ and u(x)-u(-x) solve the PDE, and thus $u(0)=0$ for all $k$ in the range. Therefore, there should be no uncertainty at the origin.
Figure 2: Inference results for the nonlinear Poisson problem. (a) Noisy data, GT solution, mean and std of inferred $u$ using B-BiLO (b) results using BPINN with $\sigma_{f}=0.1$. (c) Posterior distribution of the PDE parameter $k$ using BiLO and BPINN with different $\sigma_{f}$, compared with the reference (Metropolis-Hastings with numerical solver). (d) Accuracy-efficiency comparison: x-axis shows the -log of the std. at the orgin, larger values indicate better physical uncertainty quantification. y-axis shows the effective sample size (ESS) per second of wall time, larger values indicate more efficient sampling. As $\sigma_{f}$ decreases, BPINN becomes more accurate but less efficient. Overall B-BiLO is more accurate and efficient at uncertainty quantification.
Figure 2 compares B-BiLO and BPINN on the nonlinear Poisson problem. Panel (a) shows the noisy data, GT, mean and 2 std. region of the inferred solution $u$ from B-BiLO. The samples of $u$ satisfy the PDE accurately—evidenced by nearly zero posterior standard deviation at the origin. Panel (b) shows BPINN with $\sigma_{f}=0.1$, where the nonzero uncertainty at the origin indicates that individual samples are not solving the PDE accurately. Panel (c) presents the posterior of the PDE parameter $k$. The reference distribution is obtained using Metropolis–Hastings sampling (which is generally less efficient than HMC) combined with a accurate numerical solver. B-BiLO closely matches the reference distribution, while BPINN remains accurate only for small $\sigma_{f}$; larger $\sigma_{f}$ reduces fidelity. Panel (d) compares accuracy and efficiency. The x-axis ($\log_{10}$ of the std at origin) measures physical consistency—larger values indicate more accurate PDE solution. While the y-axis measures the efficiency using the Effective Sample Size (ESS, which quantifies the number of independent samples) per second. For each $\sigma_{f}$, we cross validated the step size $\delta t$ = 0.01, 0.005, 0.001, 0.0005 and report the best ESS/sec. As $\sigma_{f}$ decreases (color from yellow to blue), BPINN becomes more accurate but less efficient, due to the stability constraint on the step size. B-BiLO achieves both high accuracy and high efficiency simultaneously, without tuning $\sigma_{f}$. Additional visualizations are provided in SM.
Figure 3 illustrates the accuracy and efficiency of B-BiLO with LoRA of varying ranks. Panel (a) compares the posterior distribution of $k$ for LoRA ranks 4 and 8 against Full FT, demonstrating that LoRA does not compromise accuracy. Panel (b) reports the relative efficiency of LoRA (rank 8) versus Full FT, plotted as ratios with respect to Full FT. The x-axis shows the number of neurons per layer (depth fixed at 6), indicating increasing network size. The right y-axis shows the relative maximum memory usage, which remains consistently below 1, confirming LoRA’s reduced memory footprint. The left y-axis (blue line) shows the relative average wall time per sample. For small networks, LoRA is slower than Full FT (ratio above 1), while for large networks it becomes faster (ratio below 1). This reflects a trade-off: LoRA reduces the number of trainable parameters but introduces computational overhead and may require more iterations to reach comparable accuracy. As network size grows, the speedup due to reduced trainable parameters outweighs the cost of the overhead and additional iterations.
Figure 3: Accuracy and efficiency of B-BiLO with LoRA of different ranks. (a) Comparison of the posterior distribution of $k$ obtained by B-BiLO with LoRA of different ranks and Full FT. Diamond markers denote the MAP estimates, and bars indicate the 95% highest-density posterior region (HDPR). (b) Relative efficiency of LoRA (rank 8) as ratios with respect to Full FT. The x-axis shows the number of neurons per layer; the left y-axis (blue line) shows the relative average time per sample, and the right y-axis shows the relative maximum memory usage during sampling.
2.3 Example 2: Inferring Patient-Specific Tumor Growth Parameters from MRI Data
We consider a real-world application of B-BiLO for uncertainty quantification of patient-specific parameters GBM growth models using patient MRI data in 2D [zhangPersonalizedPredictionsGlioblastoma2025, balcerakIndividualizingGliomaRadiotherapy2025, ezhovLearnMorphInferNewWay2023, scheufeleFullyAutomaticCalibration2021, lipkovaPersonalizedRadiotherapyDesign2019]. Glioblastoma (GBM) is a highly aggressive brain tumor with poor prognosis, where tumor cells infiltrate beyond lesions visible on MRI, limiting the effectiveness of current treatment strategies and leading to recurrence, thereby underscoring the need for personalized prediction. The infiltration and proliferation of tumor cells can be modeled by reaction–diffusion equations [swansonQuantitativeModelDifferential2000]. Here, we use the formulation presented in [zhangPersonalizedPredictionsGlioblastoma2025]. Let $\Omega$ denote the 2D brain domain derived from MRI data, and the normalized tumor cell density $u(\mathbf{x},t)$ satisfy the Fisher–KPP equation with Neumann boundary conditions:
$\frac{\partial u}{\partial t}=D\bar{D}\gradient\cdot(P(\mathbf{x})\gradient u)+\rho\bar{\rho}u(1-u)\quad\text{in }\Omega$ (9)
where $P$ depends on the tissue distribution (e.g., white and grey matter) obtained from the MRI data, $\bar{D}$ and $\bar{\rho}$ are known patient-specific characteristic parameters based on the data, and $D$ and $\rho$ are the unknown nondimensionalized parameters that we aim to infer.
We consider two regions of interest in the tumor, the whole tumor (WT) region and the tumor core (TC) region. Let $\hat{\mathbf{y}}^{s}$, $s\in\{\rm WT,TC\}$ be indicator functions of the WT and TC regions, which can be obtained from the MRI data and serves as the observational data in the inverse problem. We define our predicted segmentations to be regions in which the tumor cell density $u$ at the nondimensional $t=1$ lies above a certain threshold $u_{c}^{s}$. The likelihood function is defined based on error between the predicted segmentation $\mathbf{y}^{s}$ and the observed segmentation $\hat{\mathbf{y}}^{s}$. We aim to find the posterior distribution of the parameters $(D,\rho,u_{c}^{\rm WT},u_{c}^{\rm TC})$. Details of the setup and validation with synthetic data are provided in the SM (see also [zhangPersonalizedPredictionsGlioblastoma2025]).
Figure 4 summarizes the inference results for one patient case. In panel (a), the posterior distributions of the parameters show that the model parameters are well identified with moderate uncertainty, and the results from B-BiLO with LoRA rank 8 closely match those from Full FT. Panel (b) shows the posterior mean and standard deviation of the tumor cell density $u$, respectively. In panel (c), the MAP segmentation is seen to closely match the observed WT and TC regions on the MRI, and the predicted infiltration margin (where $u=1\%$) extends slightly beyond the visible lesion, suggesting possible microscopic invasion. In panel (d), the standard deviation of the tumor cell density $u$ is shown. Because the data are noisy and the boundary between the TC and WT regions is poorly defined, the estimated uncertainty is higher in this transition zone. Although no direct ground truth exists for the true tumor cell density in patients, the predictive capability of the model can be indirectly validated using recurrence data [zhangPersonalizedPredictionsGlioblastoma2025, balcerakIndividualizingGliomaRadiotherapy2025, balcerakPhysicsRegularizedMultiModalImage2024]. This example illustrates the capability of B-BiLO to handle real patient data for risk-aware treatment planning; applications to full 3D patient datasets are left for future work.
Figure 4: Inference of patient-specific GBM growth parameters using B-BiLO with LoRA rank 8. (a) Posterior distribution of the parameters $D,\rho,u_{c}^{\rm WT},u_{c}^{\rm TC}$ and the prior (dashed), compared with Full FT. (b) Posterior mean of the tumor cell density $u$. (c) Ground truth MRI data with segmented WT and TC regions (filled), the MAP predicted segmentation (contour), and the predicted infiltration margin (where $u=1\%$). (d) Posterior standard deviation of the tumor cell density $u$.
2.4 Example 3: Inferring Stochastic Rates from Particle Data
We consider a problem motivated by inferring the dynamics of gene expression from static images of mRNA molecules in cells [milesInferringStochasticRates2024, milesIncorporatingSpatialDiffusion2025]. The steady-state spatial distribution of these molecules can be modeled by the following boundary value problem with a singular source:
$\frac{d^{2}u}{dx^{2}}+\lambda\delta(x-z)-\mu u=0,\quad u(0)=u(1)=0.$ (10)
Here, the solution $u(x)$ represents the intensity of a spatial Poisson point process describing the locations of mRNA particles in a simplified 1D domain, and $\delta(x-z)$ is the Dirac delta function representing a point source of mRNA at location $z$. The parameter $\lambda$ is the dimensionless birth rate (transcription) of mRNA at a specific gene site $z$, while $\mu$ is a degradation rate, and the boundary condition describes export across the nuclear boundary.
Given $M$ snapshots of mRNA locations $\{q_{i}^{j}\}$ for $j=1,\ldots,M$, $i=1,\ldots,N_{j}$, where $N_{j}$ is the number of particles in snapshot $j$, the goal is to infer the kinetic parameters $\lambda$ and $\mu$. Since the particle locations follow a Poisson point process, we can construct a likelihood function directly from their positions. The resulting log-likelihood is [milesInferringStochasticRates2024]:
$l(\lambda,\mu;q_{i}^{j})=M\int_{\Omega}u(x;\lambda,\mu)dx-\sum_{j=1}^{M}\sum_{i=1}^{N_{j}}\log u(q_{i}^{j};\lambda,\mu).$
This particular likelihood function, which is standard for spatial Poisson processes, consists of two terms: the sum of the log-intensity at each observed particle’s location and a term penalizing the total expected number of particles in the domain, $\int_{\Omega}u(x)dx$. The solution to the PDE is continuous, but its derivative is discontinuous at the source $z$; we handle this singularity by adapting the cusp-capturing PINN [tsengCuspcapturingPINNElliptic2023]. The details are explained in the SM.
In Fig.5(a), we show the 90% highest-density posterior region (HDPR) and the MAP estimates of $\lambda$ and $\mu$ obtained using the reference method, B-BiLO with LoRA (rank 4), and B-BiLO with Full FT. Both fine-tuning strategies closely match the reference, with relative errors in the MAP estimates and HDPR area below 4% (details in SM). The slight overestimation of $\lambda_{\rm MAP}$ and $\mu_{\rm MAP}$ likely arises from the inexact PDE solution, while the small discrepancies between Full FT and LoRA are attributable to variability in finite MCMC sampling. The inset shows the HDPR and MAP of the solution $u(x)$, which has a cusp at the source location $z=0.5$ and is accurately captured by B-BiLO. In Fig. 5(b), we show the ground truth (GT), prior distributions, and posterior distributions of $\lambda$ and $\mu$.
Figure 5: Inference of gene expression parameters $\lambda$ (birth rate) and $\mu$ (degradation rate) using B-BiLO. (a): The 90% highest-density posterior region (HDPR) and maximum a posteriori (MAP) estimates of the solution obtained using B-BiLO with LoRA rank 4, Full FT, and the reference method (MH with numerical solver). Inset: 90% HDPR and MAP of the solution $u(x)$. (b): Ground truth (GT), prior, and posterior distributions for $\lambda$ and $\mu$, with posterior sampled via B-BiLO using LoRA rank 4.
2.5 Example 4: Darcy Flow Problem
Our framework can be generalized to inferring unknown functions in the PDE, as described in Methods (Sec. 4.5). We consider the following 2D Darcy flow problem on the unit square $\Omega=[0,1]\times[0,1]$ with Dirichlet boundary conditions:
[…]
where $f(\mathbf{x})=10$. The unknown spatially varying diffusion coefficient $D(\mathbf{x})$ is represented by a truncated 64-term Karhunen–Loève expansion with standard normal priors on the coefficients, modeling high- and low-diffusivity regions (3 to 12) across the domain (Details are provided in the SM) [huangEfficientDerivativefreeBayesian2022].
The reference result is obtained by MH sampling with numerical solutions computed on a $\times61$ grid. Noisy observations are collected at $21\times 21$ evenly spaced grid points, and the residual loss is evaluated at $61\times 61$ collocation points. The noise level is set to $\sigma_{d}=0.01$, corresponding to approximately 10% of the maximum value of the solution $u$. Figure 6 illustrates the posterior inference of the spatially varying diffusion coefficient $D(\mathbf{x})$ in the 2D Darcy flow problem. Panel (a) shows the ground truth (GT) diffusion coefficient. Panel (b) presents the point-wise posterior means of $D(\mathbf{x})$ obtained using B-BiLO with LoRA rank 4, which captures the key spatial features of the GT. Panel (c) quantifies the uncertainty in $D(x)$ along a fixed 1D slice of the domain, indicated by the dashed line in (a), showing the region of mean $\pm$ two standard deviations. The relative L2 error of the B-BiLO posterior mean with respect to the reference is 6.2%. Additional visualizations are provided in the SM.
Figure 6: Posterior inference of the spatially varying diffusion coefficient $D(\mathbf{x})$ in the 2D Darcy flow problem. (a) The ground truth $D(\mathbf{x})$; (b) The posterior mean of $D(\mathbf{x})$ obtained using B-BiLO with LoRA rank 4; (c) Mean $\pm$ 2 standard deviations of the posterior $D(\mathbf{x})$ evaluated along a 1D slice of the domain (dashed line in (a) and (b)).
3 Discussion
In this paper, we have introduced a novel framework (B-BiLO) for efficiently solving Bayesian PDE inverse problems. Our approach enforces strong PDE constraints and integrates gradient-based MCMC methods, significantly enhancing computational efficiency and accuracy in parameter inference and uncertainty quantification. Through numerical experiments, we demonstrated that our method provides superior performance compared to Bayesian Physics-Informed Neural Networks (BPINNs). Notably, our approach avoids the computational challenges of directly sampling high-dimensional Bayesian neural network weights and mitigates issues arising from ill-conditioned PDE residual terms. Moreover, by incorporating LoRA, we substantially reduce computational overhead during the fine-tuning phase without compromising the accuracy or robustness of Bayesian inference. Our results on various problems highlight the effectiveness of our method. A promising direction for future work is extending the B-BiLO framework to three-dimensional PDEs and inverse problems with high-dimensional parameter spaces, such as full-field identification problems in geophysics or medical imaging. Tackling these challenges will likely require integration with more efficient sampling algorithms, improved neural network architectures tailored to complex PDEs, and faster fine-tuning strategies to maintain scalability.
4 Methods
4.1 Architecture
The local operator $u(\mathbf{x},\theta;W)$ is represented by a standard multi-layer perceptron (MLP). The network input is the concatenation of the coordinate and the PDE parameter, $h^{(0)}=[\mathbf{x};\theta]\in\mathbb{R}^{d+m}$. The hidden layers are defined by $h^{(l)}=\sigma(W^{(l)}h^{(l-1)}+b^{(l)})$ for $l=1,\dots,L$, where $W^{(l)}\in\mathbb{R}^{p\times d_{l-1}}$ and $b^{(l)}\in\mathbb{R}^{p}$, with $d_{0}=d+m$ and $d_{l}=p$ for all hidden layers, and $\sigma$ is the activation function ($\tanh$ in this work). The network output is $m(\mathbf{x},\theta;W)=W^{(L+1)}h^{(L)}+b^{(L+1)}$, where $W^{(L+1)}\in\mathbb{R}^{1\times p}$ and $b^{(L+1)}\in\mathbb{R}$. The collection of all trainable parameters is $W=\{W^{(l)},b^{(l)}\}_{l=0}^{L+1}$. A final transformation can be applied to the output of the MLP to enforce the boundary condition or the initial condition [dongMethodRepresentingPeriodic2021, sukumarExactImpositionBoundary2022]. For example, if the PDE is defined on a unit interval $[0,1]$ with Dirichlet boundary conditions $u(0)=u(1)=0$, the BiLO solution can be represented as $u(x,\theta;W)=m(x,\theta;W)(1-x)x$. Alternatively, the boundary condition can be imposed by additional loss terms [raissiPhysicsinformedNeuralNetworks2019]. This MLP architecture is presented as a basic example and to facilitate the explanation of Low-Rank Adaptation in Section 4.4. More advanced architectures can also be incorporated, including residual connections [heDeepResidualLearning2015], random Fourier features [wangEigenvectorBiasFourier2021], and other modified MLP designs [wangUnderstandingMitigatingGradient2021].
4.2 Algorithm
The key component of our method is to compute the gradient of the potential energy $U(\theta)$. In BiLO, the potential energy depends on the weights of the neural network and is denoted as $U(\theta,W)$. At the optimal weights $W^{*}(\theta)$, the local operator $u(\mathbf{x},\theta;W^{*}(\theta))$, the potential energy $U(\theta,W^{*}(\theta))$, and its gradient $\nabla_{\theta}U(\theta,W^{*}(\theta))$ are expected to approximate the exact local operator $u(\mathbf{x},\theta)$, the true potential energy $U(\theta)$, and the true gradient $\nabla_{\theta}U(\theta)$, respectively. Therefore, each time $\theta$ is updated, we must solve the lower level problem. The complete B-BiLO algorithm is detailed in Algorithm 1.
In practice, we solve the lower level problem to some tolerance $\epsilon$ using our subroutine LowerLevelIteration in Algorithm 2. It is written as a simple gradient descent method, but in practice, we can use more advanced optimization methods such as Adam [kingmaAdamMethodStochastic2017]. In the HMC, we call the subroutine LowerLevelIteration each time before computing the gradient of the potential energy $\nabla_{\theta}U(\theta,W)$.
Algorithm 1 B-BiLO
1:Initial states for $\theta^{(0)}$, time step size $\delta t$, and number of leapfrog steps $L$
2:for $k=1,2,\dots,N$ do
3:  Sample $r\sim\mathcal{N}(0,M)$
4:  $(\theta_{0},r_{0})\leftarrow(\theta^{(k-1)},r)$
5:  $H_{0}\leftarrow H(\theta_{0},r_{0};W)$
6:  LowerLevelIteration($\theta_{0},W$)
7:  for $i=0,1,\dots,L-1$ do
8:   $r_{i+1/2}\leftarrow r_{i}-\frac{\delta t}{2}\nabla_{\theta}U(\theta_{i},W)$
9:   $\theta_{i+1}\leftarrow\theta_{i}+\delta tM^{-1}r_{i+1/2}$
10:   LowerLevelIteration($\theta_{i+1},W$)
11:   $r_{i+1}\leftarrow r_{i+1/2}-\frac{\delta t}{2}\nabla_{\theta}U(\theta_{i+1},W)$
12:  end for
13:  $H_{L}\leftarrow H(\theta_{L},r_{L};W)$
14:  $\alpha\leftarrow\min\left(1,\exp\left(H_{0}-H_{L}\right)\right)$
15:  with probability $\alpha$, set $\theta^{(k)}\leftarrow\theta_{L}$; otherwise, set $\theta^{(k)}\leftarrow\theta_{0}$
16:end for
Algorithm 2 LowerLevelIteration($\theta$, $W$)
1:while $L_{\text{LO}}(\theta,W)>\epsilon$ do
2:  $W\leftarrow W-\alpha_{W}\nabla_{W}L_{\text{LO}}(\theta,W)$
3:end while
4.3 Theoretical analysis
The accuracy of the leapfrog integrator depends on the accuracy of the potential energy gradient, $\nabla_{\theta}U(\theta,W)$, and the accuracy of the potential posterior distribution depends on the accuracy of the PDE solution. Both errors are attributed to the inexact solution of the lower-level optimization problem.
The ideal optimal weights $W^{*}(\theta)$ satisfy the PDE for all $\theta$: $\mathcal{F}[u(\cdot,\theta,W^{*}(\theta)),\theta]=\mathbf{0}$. Its practical approximation is denoted by $\overline{W}$, which is obtained by terminating the optimization once the lower-level loss is within a specified tolerance, $\mathcal{L}_{\text{LO}}(\theta,\overline{W})\leq\epsilon$. The true gradient of the upper-level objectives is $g_{true}(\theta)=d_{\theta}U(\theta,W^{*}(\theta))$. In BiLO, we compute an approximate gradient $g_{a}(\theta)=\nabla_{\theta}U(\theta,\overline{W})$, which does not consider the dependency of $\overline{W}$ on $\theta$. Under mild assumptions of PDE operator stability, smoothness, and Lipschitz continuity of the solution map, we proved that $\norm{g_{a}-g_{true}}=O(\epsilon)$ [zhangBiLOBilevelLocal2025a].
In addition to the error of the approximate gradient, inexact lower-level optimization also introduces a static error in the posterior distribution targeted by the BiLO-HMC sampler. Specifically, denote $u^{*}=u(\cdot,\theta;W^{*}(\theta))$, and $\bar{u}=u(\cdot,\theta;\overline{W})$. The practical posterior $\bar{\pi}(\theta)$ is based on inexact solution of the PDE $\bar{u}$, while the ideal posterior $\pi_{*}(\theta)$ is based on the exact solution $u^{*}$. In Theorem 1, given in the SM, we quantify this static discrepancy by bounding the Kullback–Leibler divergence:
$D_{\mathrm{KL}}\!\left(\bar{\pi}\;\big\|\;\pi_{*}\right)=O(\epsilon).$
4.4 Efficient Sampling with Low-Rank Adaptation (LoRA)
The neural network weights $W$ are randomly initialized. Denote the initial weights as $W_{\rm rand}$. Sampling the posterior distribution $P(\theta|D)$ requires an initial guess for the PDE parameters $\theta_{0}$. During sampling, BiLO-HMC generates a sequence of PDE parameters $\theta^{(k)}_{i}$ for $k=0,1,2,\dots,N$ and $i=0,1,\dots,L$, which is the $i$-th step of the leapfrog integrator for the $k$-th proposal. The lower-level problem needs to be solved to obtain the neural network weights $W(\theta^{(k)}_{i})$. We separate the training process into two stages: pre-training and fine-tuning. And discuss how to speed up the training process.
Pre-training (Initialization)
During pre-training, the neural network weights are initialized randomly, and we fix the PDE parameter $\theta^{(0)}$ and minimize the local operator loss $\mathcal{L}_{\text{LO}}(\theta^{(0)},W)$ to obtain the initial weights $W^{*}(\theta^{(0)})$. The pre-training process can be sped up if the numerical solution of the PDE $u$ at $\theta_{0}$ is available and included as an additional data loss.
$W^{*}=\arg\min_{W}\left(\mathcal{L}_{\text{LO}}(\theta_{0},W)+\|u(\cdot,\theta_{0};W)-u(\cdot,\theta_{0})\|^{2}\right)$ (12)
The additional data loss is not mandatory for training the local operator with fixed $\theta_{0}$. However, the additional data loss can speed up the training process [krishnapriyanCharacterizingPossibleFailure2021], and is computationally inexpensive as we only need one numerical solution corresponding to $\theta_{0}$.
Fine-tuning (Sampling)
In the fine-tuning stage, we train the local operator for the sequence $\theta^{(k)}_{i}$ for $k=0,1,2,\dots,N$ and $i=0,1,\dots,L$, where the changes in the PDE parameters are small. We do not train from scratch. Instead, we fine-tune the neural network weights $W$ based on the previous weights, which usually converges much faster.
In Algorithm 2, we fine-tune all the weights of the neural network $W$. We call this approach full fine-tuning (Full FT). For large neural networks, this can still be computationally expensive. In the next section, we describe how to use the Low-Rank Adaptation (LoRA) technique [huLoRALowRankAdaptation2021] to reduce the number of trainable parameters and speed up the fine-tuning process.
Efficient Fine-tuning with Low Rank Adaptation (LoRA)
Low-Rank Adaptation (LoRA) is a technique initially developed for large language models [huLoRALowRankAdaptation2021]. In LoRA, instead of updating the full weight matrix, we learn an low-rank modification to the weight matrix. Below we describe the details of LoRA in the context of BiLO, following the notations used in Sec. 4.1.
Suppose the initial guess of the PDE parameter is $\theta^{(0)}$. After pre-training, let ${\mathbf{W}}^{(l)}_{0}\in\mathbb{R}^{p\times p}$ be the weight matrix of the $l$-th layer of the neural network after pre-training, and $W_{0}$ denote the collection of all weight matrices of the neural network after pre-training at fixed $\theta^{(0)}$.
During sampling, after each step of the leapfrog integrator, we arrive at some different PDE parameters $\theta$. We need to solve the lower level problem again to obtain the new weight matrix ${\mathbf{W}^{(l)}}(\theta)$. Consider the effective update of the weight matrix
$\Delta\mathbf{W}^{(l)}(\theta)=\mathbf{W}^{(l)}(\theta)-\mathbf{W}^{(l)}_{0}$.
In LoRA, it is assumed that the change in weights has a low-rank structure, which can be expressed as a product of two low-rank matrices:
$\Delta\mathbf{W}^{(l)}=\mathbf{A}^{(l)}\mathbf{B}^{(l)},$
where $\mathbf{A}^{(l)}\in\mathbb{R}^{p\times r}$ and $\mathbf{B}^{(l)}\in\mathbb{R}^{r\times p}$ are low-rank matrices with $r\ll p$.
Denote the collection of low-rank matrices as $W_{LoRA}=\{\mathbf{A}^{(l)},\mathbf{B}^{(l)}\}_{l=1}^{L}$. Then neural network is represented as $u(\mathbf{x},\theta;W_{0}+W_{LoRA})$, where $W_{0}+W_{LoRA}$ means that the $l$-th layer weight matrix is given by $\mathbf{W}^{(l)}=\mathbf{W}^{(l)}_{0}+\mathbf{A}^{(l)}\mathbf{B}^{(l)}$. This approach aims to reduce the complexity of the model while preserving its ability to learn or adapt effectively. With LoRA, our lower level problem becomes
$W_{LoRA}^{*}(\theta)=\arg\min_{W_{LoRA}}\mathcal{L}_{\text{LO}}(\theta,W_{0}+W_{LoRA})$ (13)
where $\theta$ are the PDE parameters computed in leapfrog. The full FT subroutine (Algorithm 2) can be replaced by the LoRA fine-tuning in Algorithm 3. In this work, we use the original LoRA [huLoRALowRankAdaptation2021] as a proof-of-concept and we note that more recent developments of LoRA might further increase efficiency and performance [zhangAdaLoRAAdaptiveBudget2023, wangLoRAGALowRankAdaptation2024, hayouLoRAEfficientLow2024].
Algorithm 3 LowerLevelIteration-LoRA($\theta$)
1:Initial weights $W_{0}$ after pre-training, rank $r$, learning rate $\alpha_{W_{LoRA}}$, and stopping tolerance $\epsilon$
2:while $L_{\text{LO}}(\theta,W_{0}+W_{LoRA})>\epsilon$ do
3:  $W_{LoRA}\leftarrow W_{LoRA}-\alpha_{W_{LoRA}}\nabla_{W_{LoRA}}\mathcal{L}_{\text{LO}}(\theta,W_{0}+W_{LoRA})$
4:end while
5:return $W_{0}+W_{LoRA}$
Trade-off of LoRA
The number of trainable weights $W_{LoRA}$ is much smaller than the full weight matrix $W$, and therefore LoRA reduces the memory requirements significantly at each iteration. However, evaluating the neural network (forward pass) and computing the gradient via back propagation still requires the full weight matrix $W=W_{0}+W_{LoRA}$. Using the additional low-rank matrices $\mathbf{A}^{(l)}$ and $\mathbf{B}^{(l)}$ might incur additional computational overhead. Thus, LoRA is more beneficial for large neural networks (in our case, large $p$).
In addition, with LoRA, the optimization problem is solved in a lower dimensional subspace of the full weight space. Therefore, it is expected that more iterations are needed to reach the same level of loss as full-model fine-tuning [chenCELoRAComputationEfficientLoRA2025]. On the other hand, optimizing over a subspace can serve as a regularization [bidermanLoRALearnsLess2024]. In numerical experiments, we show that as the model size increases, the benefit outweighs the cost, and LoRA can significantly speed up the fine-tuning process.
4.5 Inferring Unknown Functions
We describe in general how to infer unknown functions in the BiLO framework [zhangBiLOBilevelLocal2025a]. Suppose the PDE depends on some unknown functions $f(\mathbf{x})$, such as the spatially varying diffusivity in a diffusion equation. Denote the PDE as
$F(D^{k}u(\mathbf{x}),...,Du(\mathbf{x}),u(\mathbf{x}),f(\mathbf{x}))=\mathbf{0}$ (14)
We introduce an auxiliary variable $z=f(\mathbf{x})$, and we aim to find a local operator $u(\mathbf{x},z)$ such that $u(\mathbf{x},f(\mathbf{x}))$ solves the PDE locally at $f$. We introduce the augmented residual function, which has an auxiliary variable $z$:
$a(\mathbf{x},z):=F(D^{k}u(\mathbf{x},z),...,Du(\mathbf{x},z),u(\mathbf{x},z),z).$ (15)
And the conditions for a local operator are:
  1. 1.
    $a(\mathbf{x},f(\mathbf{x}))=\mathbf{0}$
  2. 2.
    $\gradient_{z}a(\mathbf{x},z)|_{z=f(\mathbf{x})}=d_{z}\mathcal{F}(D^{k}u(\mathbf{x},z),...,Du(\mathbf{x},z),u(\mathbf{x},z),z)|_{z=f(\mathbf{x})}=\mathbf{0}$.
Condition 1 states that the function $u(\mathbf{x},f(\mathbf{x}))$ has zero residual, and condition 2 means that small perturbations of $f(\mathbf{x})$ result in small changes in the residual, e.g., the PDE is approximated to second order in the perturbation size (see also [zhangBiLOBilevelLocal2025a]).
The unknown function $f$ can be represented by a Bayesian neural network [nealBayesianLearningNeural1996, yangBPINNsBayesianPhysicsinformed2021] or a Karhunen-Loève (KL) expansion [rasmussenGaussianProcessesMachine, huangEfficientDerivativefreeBayesian2022, liFourierNeuralOperator2024]. We denote the parameterized function by $f(\mathbf{x},\theta)$, where $\theta$ are the weights of the Bayesian neural network or the coefficients of the KL expansion. The Local Operator loss in this case is
$\mathcal{L}_{\text{LO}}(\theta,W)=\norm{a(\cdot,f(\cdot,\theta),W)}^{2}_{2}+w_{\rm rgrad}\norm{\nabla_{z}a(\cdot,f(\cdot,\theta),W)}^{2}_{2},$ (16)
and the same bi-level Bayesian inference problem (5) is solved.
[…]
Code Availability The code for the numerical experiments is available at https://github.com/Rayzhangzirui/BiLO.
[…]
Acknowledgments R.Z.Z and J.S.L thank Babak Shahbaba for the GPU resources. R.Z.Z. thanks the NVIDIA Academic Grant Program for the NVIDIA RTX PRO 6000 Blackwell GPU. J.S.L acknowledges partial support from the National Science Foundation through grants DMS-2309800, DMS-1953410 and DMS-1763272 and the Simons Foundation (594598QN) for an NSF-Simons Center for Multiscale Cell Fate Research. C.E.M was partially supported by a NSF CAREER grant DMS-2339241.